# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

### v1

@@ Abstract.
Our contributions are: (i) pushouts as higher-inductive types with explicit path constructors; (ii) free products and amalgamated free products as quotients of word representations; (iii) an encode-decode proof establishing $\pi_{1}(\mathrm{Pushout}(A,B,C))\simeq\pi_{1}(A)\ast_{\pi_{1}(C)}\pi_{1}(B)$; and (iv) applications to the figure-eight ($\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$) and 2-sphere ($\pi_{1}(S^{2})\simeq 1$).
The framework makes coherence witnesses explicit as rewrite derivations. The development is formalized in Lean 4: pushout axioms and the encode map are assumed, while the decode map, amalgamation compatibility, and applications are fully mechanized ($\sim$2050 lines). This demonstrates that the encode-decode method for higher-inductive types becomes fully constructive when path equality is decidable via normalization.
@@ 1. Introduction
The Seifert-van Kampen theorem, first proved by Herbert Seifert [15] and Egbert van Kampen [17], is one of the most powerful tools in algebraic topology for computing fundamental groups (see [8] for a modern treatment). It states that if a path-connected space $X$ is the union of two path-connected open subspaces $U$ and $V$ with path-connected intersection $U\cap V$, then the fundamental group $\pi_{1}(X)$ can be computed as the amalgamated free product:
@@ 1.3. Structure of the Paper
Section 2 reviews computational paths, rewrite equality, and fundamental groups. Section 3 defines pushouts as higher-inductive types. Section 4 constructs free products and amalgamated free products. Section 5 proves the Seifert-van Kampen theorem using the encode-decode method. Section 6 applies the theorem to compute fundamental groups of the figure-eight and 2-sphere. Section 7 discusses the Lean 4 formalization. Section 8 concludes.
@@ 2.1. Computational Paths
  • Transitivity: If $p:\mathrm{Path}(a,b)$ and $q:\mathrm{Path}(b,c)$, then $\tau(p,q):\mathrm{Path}(a,c)$ — concatenate the sequences.
  • Congruence: If $p:\mathrm{Path}(a,b)$ and $f:A\to B$, then $\mathrm{congrArg}(f,p):\mathrm{Path}(f(a),f(b))$.
@@
In Lean, a path is represented as:
structure Path {A : Type u} (a b : A) where
toEq : a = b -- the underlying equality proof
The key insight is that we distinguish paths by their structure, not just by the fact that they connect the same endpoints.
@@ 2.2. The Step Relation
  • Groupoid laws: $\sigma(\rho)\rightsquigarrow\rho$, $\sigma(\sigma(p))\rightsquigarrow p$, $\tau(\rho,p)\rightsquigarrow p$, $\tau(p,\rho)\rightsquigarrow p$, $\tau(p,\sigma(p))\rightsquigarrow\rho$, $\tau(\sigma(p),p)\rightsquigarrow\rho$, $\sigma(\tau(p,q))\rightsquigarrow\tau(\sigma(q),\sigma(p))$, and $\tau(\tau(p,q),r)\rightsquigarrow\tau(p,\tau(q,r))$.
@@
inductive Step : {A : Type u} -> {a b : A} -> Path a b -> Path a b -> Prop
| symm_refl (a : A) : Step (symm (Path.refl a)) (Path.refl a)
| symm_symm (p : Path a b) : Step (symm (symm p)) p
| trans_refl_left (p : Path a b) : Step (trans (Path.refl a) p) p
| trans_refl_right (p : Path a b) : Step (trans p (Path.refl b)) p
| trans_symm (p : Path a b) : Step (trans p (symm p)) (Path.refl a)
| symm_trans (p : Path a b) : Step (trans (symm p) p) (Path.refl b)
| trans_assoc (p : Path a b) (q : Path b c) (r : Path c d) :
Step (trans (trans p q) r) (trans p (trans q r))
-- ... plus type-specific and context rules
@@ 2.3. Rewrite Equality
$\tau(\rho(a),p)$ $\rightsquigarrow p$ (left unit)
$\tau(p,\rho(b))$ $\rightsquigarrow p$ (right unit)
$\tau(p,\sigma(p))$ $\rightsquigarrow\rho(a)$ (right inverse)
$\tau(\sigma(p),p)$ $\rightsquigarrow\rho(b)$ (left inverse)
$\tau(\tau(p,q),r)$ $\rightsquigarrow\tau(p,\tau(q,r))$ (associativity)
@@
inductive RwEq {A : Type u} {a b : A} : Path a b -> Path a b -> Prop
| refl (p : Path a b) : RwEq p p
| step {p q : Path a b} : Step p q -> RwEq p q
| symm {p q : Path a b} : RwEq p q -> RwEq q p
| trans {p q r : Path a b} : RwEq p q -> RwEq q r -> RwEq p r
@@
    Completeness with respect to canonical forms: Any path $p$ is rewrite-equivalent to its normal form $\mathrm{nf}(p)$, so $p\sim\mathrm{nf}(p)$.
@@
    Decidability: $\sim$ is decidable by comparing normal forms: $p\sim q$ iff $\mathrm{nf}(p)=\mathrm{nf}(q)$.
Soundness is immediate from the definition of $\sim$—each rewrite rule preserves the underlying equality. The completeness property follows from confluence: any path can be reduced to its unique normal form. In our Lean formalization, we additionally have that paths between the same endpoints with the same underlying equality proof are provably $\sim$, since Lean’s definitional equality on the Path structure respects the stored equality proof. Together these yield:
Theorem 1 (Canonicity).
For any paths $p,q:\mathrm{Path}(a,b)$, we have $p\sim q$ if and only if $p.\mathrm{toEq}=q.\mathrm{toEq}$.
This theorem is the key that allows $\pi_{1}$ to be computed effectively: path equality reduces to equality of underlying proofs, which in turn reduces to normalization.
@@ 2.4. Loop Spaces and Fundamental Groups
{defi}
[Loop Space] The loop space of a type $A$ at a basepoint $a:A$ is:
@@
{defi}
[Fundamental Group] The fundamental group $\pi_{1}(A,a)$ is the quotient of the loop space by rewrite equality:
@@
abbrev LoopSpace (A : Type u) (a : A) : Type u := Path a a
def PiOne (A : Type u) (a : A) : Type u := Quot (@RwEq A a a)
notation "pi_1(" A "," a ")" => PiOne A a
@@
  • Identity: $\rho(a)]$
  • Multiplication: $\alpha]\cdot[\beta]:=[\tau(\alpha,\beta)]$
  • Inverse: $\alpha]^{-1}:=[\sigma(\alpha)]$
@@ 3.1. The Pushout Type
{defi}
[Pushout] The pushout $\mathrm{Pushout}(A,B,C,f,g)$ is characterized by:
@@
axiom Pushout (A : Type u) (B : Type u) (C : Type u)
(f : C -> A) (g : C -> B) : Type u
namespace Pushout
axiom inl : A -> Pushout A B C f g
axiom inr : B -> Pushout A B C f g
axiom glue (c : C) : Path (inl (f c)) (inr (g c))
end Pushout
@@ 3.2. Recursion Principle
{defi}
[Pushout Recursion] Given a type $D$ with:
@@
$\mathrm{congrArg}(\mathrm{rec},\mathrm{glue}(c))$ $\sim\mathrm{onGlue}(c)$
@@ 3.3. Glue Naturality
Lemma 2 (Glue Naturality).
@@
$\mathrm{inl}(f(c_{1}))$$\mathrm{inl}(f(c_{2}))$$\mathrm{inr}(g(c_{1}))$$\mathrm{inr}(g(c_{2}))$$\scriptstyle{\mathrm{inlPath}(f_{*}(p))}$$\scriptstyle{\mathrm{glue}(c_{1})}$$\scriptstyle{\mathrm{glue}(c_{2})}$$\scriptstyle{\mathrm{inrPath}(g_{*}(p))}$
@@
$\tau(\mathrm{inlPath}(f_{*}(p)),\mathrm{glue}(c_{2}))\sim\tau(\mathrm{glue}(c_{1}),\mathrm{inrPath}(g_{*}(p)))$
where $f_{*}(p)=\mathrm{congrArg}(f,p)$ and $g_{*}(p)=\mathrm{congrArg}(g,p)$.
@@
axiom glue_natural {c1 c2 : C} (p : Path c1 c2) :
Path.trans (inlPath (Path.congrArg f p)) (glue c2) =
Path.trans (glue c1) (inrPath (Path.congrArg g p))
axiom glue_natural_rweq {c1 c2 : C} (p : Path c1 c2) :
RwEq (inlPath (Path.congrArg f p))
(Path.trans (glue c1)
(Path.trans (inrPath (Path.congrArg g p))
(Path.symm (glue c2))))
The second form states that a path in $\mathrm{inl}(A)$ is $\sim$-equivalent to a “zig-zag” path that crosses to $\mathrm{inr}(B)$ and back.
@@ 3.4.1. Wedge Sum
$A\vee B:=\mathrm{Pushout}(A,B,\mathrm{Unit},\lambda\_.a_{0},\lambda\_.b_{0})$
@@
def Wedge (A : Type u) (B : Type u) (a0 : A) (b0 : B) : Type u :=
Pushout A B PUnit (fun _ => a0) (fun _ => b0)
@@ 3.4.2. Suspension
The suspension $\Sigma A$ is the pushout of $\mathrm{pt}\leftarrow A\to\mathrm{pt}$:
$\Sigma A:=\mathrm{Pushout}(\mathrm{Unit},\mathrm{Unit},A,\lambda\_.*,\lambda\_.*)$
@@
def Suspension (A : Type u) : Type u :=
Pushout PUnit PUnit A (fun _ => PUnit’.unit) (fun _ => PUnit’.unit)
noncomputable def north : Suspension A := Pushout.inl PUnit’.unit
noncomputable def south : Suspension A := Pushout.inr PUnit’.unit
noncomputable def merid (a : A) : Path north south := Pushout.glue a

### v2

@@ Abstract.
Our contributions are: (i) pushouts as higher-inductive types with explicit path constructors; (ii) free products and amalgamated free products as quotients of word representations; (iii) an encode-decode proof establishing $\pi_{1}(\mathrm{Pushout}(A,B,C))\simeq\pi_{1}(A)\ast_{\pi_{1}(C)}\pi_{1}(B)$; and (iv) applications to the figure-eight ($\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$), 2-sphere ($\pi_{1}(S^{2})\simeq 1$), and 3-sphere ($\pi_{1}(S^{3})\simeq 1$ with Hopf fibration context).
Recent extensions include: higher homotopy groups $\pi_{n}$ via the weak $\omega$-groupoid structure (with $\pi_{2}$ abelian via Eckmann-Hilton, and $\pi_{2}\cong 1$ in the 1-groupoid truncated setting); truncation levels connecting the framework to HoTT n-types; automated path simplification tactics; basic covering space theory with $\pi_{1}$-actions on fibers; fibration theory with long exact sequences; and Eilenberg-MacLane space characterization ($S^{1}=K(\mathbb{Z},1)$).
The framework makes coherence witnesses explicit as rewrite derivations. The development is formalized in Lean 4 with over 25,000 lines of mechanized proofs. This demonstrates that the encode-decode method for higher-inductive types becomes fully constructive when path equality is decidable via normalization.
@@ 1. Introduction
The Seifert-van Kampen theorem, first proved by Herbert Seifert [16] and Egbert van Kampen [18], is one of the most powerful tools in algebraic topology for computing fundamental groups (see [7, 11] for modern treatments). It states that if a path-connected space $X$ is the union of two path-connected open subspaces $U$ and $V$ with path-connected intersection $U\cap V$, then the fundamental group $\pi_{1}(X)$ can be computed as the amalgamated free product:
@@
Relation to Classical SVK.
The classical theorem requires $X=U\cup V$ with $U,V,U\cap V$ open and path-connected. The HIT version generalizes this: the pushout $\mathrm{Pushout}(A,B,C,f,g)$ corresponds to the union where $f$ and $g$ specify how the “intersection” $C$ embeds into both spaces. Path-connectedness assumptions are needed to ensure the fundamental group is well-defined; in the HIT setting, this becomes the assumption that all types are path-connected.
@@ 1.1. Main Contributions
    • The 3-sphere: $\pi_{1}(S^{3})\simeq 1$ with Hopf fibration context
@@ 1.2. Related Work
1.3. Reading Guide
Readers interested in the mathematical content can focus on Sections 36, treating Lean code as pseudocode. Those interested in formalization should additionally consult Section 7 and the supplementary code. Key definitions are boxed for quick reference. We use the figure-eight space $S^{1}\vee S^{1}$ as a running example throughout.
@@ 1.4. Structure of the Paper
Section 2 reviews computational paths, rewrite equality, and fundamental groups. Section 3 defines pushouts as higher-inductive types. Section 4 constructs free products and amalgamated free products. Section 5 proves the Seifert-van Kampen theorem using the encode-decode method. Section 6 applies the theorem to compute fundamental groups of the figure-eight, 2-sphere, and 3-sphere (with Hopf fibration context). Section 7 discusses the Lean 4 formalization. Section 8 concludes.
@@ 2.1. Computational Paths
  • Transitivity: If $p:\mathrm{Path}(a,b)$ and $q:\mathrm{Path}(b,c)$, then $p\cdot q:\mathrm{Path}(a,c)$ — concatenate the sequences.
  • Congruence: If $p:\mathrm{Path}(a,b)$ and $f:A\to B$, then $f_{*}(p):\mathrm{Path}(f(a),f(b))$.
@@
Notation.
Throughout this paper, we use:
  • $p\cdot q$ for path composition (transitivity)
  • $p^{-1}$ for path inverse (symmetry)
  • $f_{*}(p)$ for $\mathrm{congrArg}(f,p)$
  • $\rho(a)$ or simply $\rho$ for reflexivity
In Lean, a path is represented as a wrapper around an equality proof:
1structure Path {A : Type u} (a b : A) where
2 toEq : a = b -- the underlying equality proof
Remark 2.1 (Implementation Note).
The Lean implementation stores only the underlying equality proof, not an explicit rewrite sequence. The “rewrite sequence” interpretation is recovered via the Step and RwEq relations, which classify paths syntactically. This design choice simplifies the formalization while preserving the key property that path equality is captured by rewrite equivalence.
@@ 2.2. The Step Relation
  • Groupoid laws: $\rho^{-1}\rightsquigarrow\rho$, $(p^{-1})^{-1}\rightsquigarrow p$, $\rho\cdot p\rightsquigarrow p$, $\cdot\rho\rightsquigarrow p, pp$$\cdot p^{-1}\rightsquigarrow\rho$, $p^{-1}\cdot p\rightsquigarrow\rho$, $(p\cdot q)^{-1}\rightsquigarrow q^{-1}\cdot p^{-1}$, and $(p\cdot q)\cdot r\rightsquigarrow p\cdot(q\cdot r)$.
@@
  • Canonicalization: $p\rightsquigarrow\mathrm{ofEq}(p.\mathrm{toEq})$ — any path can reduce to its canonical form.
Remark 2.2 (Critical Pairs).
The confluence and termination of $\mathrm{LND}_{\mathrm{EQ}}$-TRS are established in [12] via analysis of 56 critical pairs. The key insight is that the canon rule—which rewrites any path $p$ to $\mathrm{ofEq}(p.\mathrm{toEq})$—acts as a “normalization shortcut” that collapses all parallel paths to a common canonical form.
@@
1inductive Step : {A : Type u} -> {a b : A} -> Path a b -> Path a b -> Prop
2 | symm_refl (a : A) : Step (symm (Path.refl a)) (Path.refl a)
3 | symm_symm (p : Path a b) : Step (symm (symm p)) p
4 | trans_refl_left (p : Path a b) : Step (trans (Path.refl a) p) p
5 | trans_refl_right (p : Path a b) : Step (trans p (Path.refl b)) p
6 | trans_symm (p : Path a b) : Step (trans p (symm p)) (Path.refl a)
7 | symm_trans (p : Path a b) : Step (trans (symm p) p) (Path.refl b)
8 | trans_assoc (p : Path a b) (q : Path b c) (r : Path c d) :
9 Step (trans (trans p q) r) (trans p (trans q r))
10 | canon (p : Path a b) : Step p (Path.ofEq p.toEq)
11 -- ... plus type-specific and context rules (50+ total)
@@ 2.3. Rewrite Equality
$\rho\cdot p$ $\rightsquigarrow p$ (left unit)
$p\cdot\rho$ $\rightsquigarrow p$ (right unit)
$p\cdot p^{-1}$ $\rightsquigarrow\rho$ (right inverse)
$p^{-1}\cdot p$ $\rightsquigarrow\rho$ (left inverse)
$(p\cdot q)\cdot r$ $\rightsquigarrow p\cdot(q\cdot r)$ (associativity)
Example 2.3 (Rewrite Derivation).
Consider $p\cdot(q\cdot q^{-1})$ where $p,q:\mathrm{Path}(a,a)$. The derivation to $p$ proceeds:
$p\cdot(q\cdot q^{-1})$ $\rightsquigarrow p\cdot\rho$ (right inverse)
$\rightsquigarrow p$ (right unit)
This explicit derivation is the “computational content” witnessing that $p\cdot(q\cdot q^{-1})\sim p$.
@@
1inductive RwEq {A : Type u} {a b : A} : Path a b -> Path a b -> Prop
2 | refl (p : Path a b) : RwEq p p
3 | step {p q : Path a b} : Step p q -> RwEq p q
4 | symm {p q : Path a b} : RwEq p q -> RwEq q p
5 | trans {p q r : Path a b} : RwEq p q -> RwEq q r -> RwEq p r
@@
    Completeness: Any path $p$ is rewrite-equivalent to its canonical form, so $p\sim\mathrm{ofEq}(p.\mathrm{toEq})$.
@@
    Decidability: $\sim$ is decidable by comparing canonical forms: $p\sim q$ iff their canonical forms are identical.
Theorem 2.4 (Canonicity).
In Lean 4 (with proof-irrelevant Prop), for any paths $p,q:\mathrm{Path}(a,b)$, we have $p\sim q$.
Proof.
Since $\mathrm{Path}(a,b)$ stores a proof $p.\mathrm{toEq}:a=b$ in Lean’s proof-irrelevant $\mathsf{Prop}$, we have $p.\mathrm{toEq}=q.\mathrm{toEq}$ definitionally. The canon rule gives $p\sim\mathrm{ofEq}(p.\mathrm{toEq})$ and $q\sim\mathrm{ofEq}(q.\mathrm{toEq})$. Since $p.\mathrm{toEq}=q.\mathrm{toEq}$, these canonical forms are identical, so $p\sim q$ by transitivity. ∎
Remark 2.5 ($\sim$ Triviality in Lean).
In our Lean 4 implementation, since Path.toEq stores a proof in the proof-irrelevant Prop type, any two paths $p,q:\mathrm{Path}(a,b)$ satisfy $p.\mathrm{toEq}=q.\mathrm{toEq}$ definitionally. Thus $\sim$ is trivially satisfied for all parallel paths. The groupoid laws become derivable rather than requiring explicit rewrite sequences.
This differs from the original computational paths theory where paths carry explicit rewrite traces. The advantage is simpler formalization; the tradeoff is that we lose the ability to distinguish paths by their internal structure. Future work could implement paths as syntactic terms (not carrying equality proofs) to recover the intended distinction.
@@ 2.4. Loop Spaces and Fundamental Groups
Definition 2.6 (Path-Connected).
A type $A$ is path-connected if for all $a,b:A$, there exists $p:\mathrm{Path}(a,b)$. Equivalently, $\prod_{a,b:A}\|\mathrm{Path}(a,b)\|_{-1}$ where $\|-\|_{-1}$ denotes propositional truncation.
Definition 2.7 (Loop Space).
The loop space of a type $A$ at a basepoint $a:A$ is:
@@
Definition 2.8 (Fundamental Group).
The fundamental group $\pi_{1}(A,a)$ is the quotient of the loop space by rewrite equality:
@@
1abbrev LoopSpace (A : Type u) (a : A) : Type u := Path a a
2
3def PiOne (A : Type u) (a : A) : Type u := Quot (@RwEq A a a)
4
5notation "pi_1(" A "," a ")" => PiOne A a
@@
  • Identity: $\rho]$
  • Multiplication: $\alpha]\cdot[\beta]:=[\alpha\cdot\beta]$
  • Inverse: $\alpha]^{-1}:=[\alpha^{-1}]$
@@
Remark 2.9 (Base Point Dependence).
The fundamental group $\pi_{1}(A,a)$ depends on the choice of basepoint $a$. For different basepoints $a,b:A$ in a path-connected type, the fundamental groups are isomorphic via conjugation by a path $\gamma:\mathrm{Path}(a,b)$:
$\pi_{1}(A,a)\cong\pi_{1}(A,b),\quad[\alpha]\mapsto[\gamma^{-1}\cdot\alpha\cdot\gamma]$
However, this isomorphism is not canonical—it depends on the choice of $\gamma$.
@@ 3.1. The Pushout Type
Definition 3.1 (Pushout).
The pushout $\mathrm{Pushout}(A,B,C,f,g)$ is characterized by:
@@
1axiom Pushout (A : Type u) (B : Type u) (C : Type u)
2 (f : C -> A) (g : C -> B) : Type u
3
4namespace Pushout
5 axiom inl : A -> Pushout A B C f g
6 axiom inr : B -> Pushout A B C f g
7 axiom glue (c : C) : Path (inl (f c)) (inr (g c))
8end Pushout
@@ 3.2. Recursion Principle
Definition 3.2 (Pushout Recursion).
Given a type $D$ with:
@@
$(\mathrm{rec})_{*}(\mathrm{glue}(c))$ $\sim\mathrm{onGlue}(c)$
@@ 3.3. Glue Naturality
Lemma 3.3 (Glue Naturality).
@@
$\mathrm{inl}(f(c_{1}))$$\mathrm{inl}(f(c_{2}))$$\mathrm{inr}(g(c_{1}))$$\mathrm{inr}(g(c_{2}))$$\scriptstyle{\mathrm{inl}_{*}(f_{*}(p))}$$\scriptstyle{\mathrm{glue}(c_{1})}$$\scriptstyle{\mathrm{glue}(c_{2})}$$\scriptstyle{\mathrm{inr}_{*}(g_{*}(p))}$
@@
$\mathrm{inl}_{*}(f_{*}(p))\cdot\mathrm{glue}(c_{2})\sim\mathrm{glue}(c_{1})\cdot\mathrm{inr}_{*}(g_{*}(p))$
Rearranging, we obtain a form useful for the decode function:
Corollary 3.4.
For any path $p:\mathrm{Path}(c_{1},c_{2})$ in $C$:
$\mathrm{inl}_{*}(f_{*}(p))\sim\mathrm{glue}(c_{1})\cdot\mathrm{inr}_{*}(g_{*}(p))\cdot\mathrm{glue}(c_{2})^{-1}$
@@
1axiom glue_natural {c1 c2 : C} (p : Path c1 c2) :
2 Path.trans (inlPath (Path.congrArg f p)) (glue c2) =
3 Path.trans (glue c1) (inrPath (Path.congrArg g p))
4
5axiom glue_natural_rweq {c1 c2 : C} (p : Path c1 c2) :
6 RwEq (inlPath (Path.congrArg f p))
7 (Path.trans (glue c1)
8 (Path.trans (inrPath (Path.congrArg g p))
9 (Path.symm (glue c2))))
Remark 3.5 (Two Forms of Glue Naturality).
The axiom glue_natural provides a propositional equality of paths, while glue_natural_rweq provides the $\sim$ equivalence needed for quotient arguments. In a cubical type theory, both would be provable; here, we axiomatize them.
@@ 3.4.1. Wedge Sum
$A\vee B:=\mathrm{Pushout}(A,B,\mathbf{1},\lambda\_.a_{0},\lambda\_.b_{0})$
@@
Example 3.6 (Figure-Eight as Wedge).
The figure-eight space is $S^{1}\vee S^{1}$—two circles joined at a single point. The glue path identifies the basepoints of the two circles:
$\mathrm{glue}(*):\mathrm{Path}(\mathrm{inl}(\mathrm{base}_{1}),\mathrm{inr}(\mathrm{base}_{2}))$
@@
1def Wedge (A : Type u) (B : Type u) (a0 : A) (b0 : B) : Type u :=
2 Pushout A B PUnit (fun _ => a0) (fun _ => b0)
@@ 3.4.2. Suspension
The suspension $\Sigma A$ is the pushout of $\mathbf{1}\leftarrow A\to\mathbf{1}$:
$\Sigma A:=\mathrm{Pushout}(\mathbf{1},\mathbf{1},A,\lambda\_.*,\lambda\_.*)$
@@
1def Suspension (A : Type u) : Type u :=
2 Pushout PUnit PUnit A (fun _ => PUnit’.unit) (fun _ => PUnit’.unit)
3
4noncomputable def north : Suspension A := Pushout.inl PUnit’.unit
5noncomputable def south : Suspension A := Pushout.inr PUnit’.unit
6noncomputable def merid (a : A) : Path north south := Pushout.glue a